# GravSphere2: A higher-order Jeans method for mass-modeling spherical stellar systems

### v1

@@ 2 Modeling stellar kinematics with GravSphere2
is the ‘velocity anisotropy’, [^7] Unlike in e.g. Bañares-Hernández et al. (2025), $\sigma_{t}$ corresponds to the total 2D tangential velocity dispersion, rather than the 1D averaged one (thus the factor of two). This is purely notational, but we have adopted this change because we found it more convenient to work with this quantity for the subsequent parts of this section. and we assume Newtonian weak-field gravity with an enclosed mass $M(r)$ within the 3D radial coordinate $r$. The solution for the radial velocity dispersion is then given by (Mamon & Łokas, 2005; van der Marel, 1994):
@@
By considering these three separate components, provided that one has enough data, it is possible to alleviate the well-known problem of the degeneracy between the cumulative mass, $M(r)$, and stellar velocity anisotropy, $\beta(r)$ (Merrifield & Kent, 1990; Wilkinson et al., 2002; Łokas & Mamon, 2003; de Lorenzi et al., 2009; Read & Steger, 2017). Further, one can resort to higher moments of the velocity distribution to place additional constraints that have also been found to alleviate this problem. One popular choice is to use VSPs (Merrifield & Kent, 1990; Richardson & Fairbairn, 2014; Read & Steger, 2017), which are two numbers obtained by integrating the luminosity-weighted, fourth-order LOS velocity moment to infinity, thereby marginalizing out any higher-order velocity anisotropy. This approach is used in GravSphere1 and GravSphere1.5.
The key advantage of using VSPs over explicit modeling of the higher-order moment equations is that no higher-order analog of $\beta(r)$ is required. However, a key disadvantage is that the VSPs require integrals to infinity over the surface brightness profile and fourth velocity moment, which may both be poorly constrained at large radii. Furthermore, their large radius behavior could be adversely affected by disequilibrium due to tides. Both effects can lead to bias (e.g. Read & Steger, 2017; Read et al., 2018; De Leo et al., 2024; Tchiorniy & Genina, 2025). For this reason, in GravSphere2 we take a different approach that – as we shall show (Sec. 4) – is less biased. Instead of using VSPs, we explicitly solve the fourth-order Jeans equations, introducing a flexible functional form for the associated higher-order anisotropy that we can then marginalize over. This is more similar to the approach introduced in Richardson & Fairbairn (2013). Explicitly, we solve the fourth-order moment Jeans equations (Merrifield & Kent, 1990):
@@
As shown in Richardson & Fairbairn (2013), this system of equations can be further simplified, without loss of generality, by first introducing the fourth-order anisotropy analog:
@@
As for the second-order counterparts, we now seek to obtain the projected quantities which are directly related to observations. To express these purely in terms of $\braket{v^{4}_{r}}$, following Richardson & Fairbairn (2013), first, it follows from Eq. (12) that:
@@
As noted in Richardson & Fairbairn (2013), this approach constitutes a significant theoretical improvement over second-order Jeans methods by ensuring physical consistency of a given mass model with higher-order velocity moments. This is not trivially warranted, as Jeans models can in some cases lead to unphysical solutions implying negative distribution functions, or be inconsistent with the velocity distribution that is implicitly assumed to map velocity dispersions (Dejonghe & Merritt, 1992; Richardson & Fairbairn, 2013) (this is explicitly addressed in the Appendix A). It also has the promise of serving as a means to further constrain mass models by considering additional higher-order moments (Richardson & Fairbairn, 2013).
One potential disadvantage, however, which is also noted in Richardson & Fairbairn (2013), is that the constraining power of this approach can be compromised due to the fact that a fully general treatment leads to the introduction of the fourth-order anisotropy analog $\beta^{\prime},$ meaning that one may not be able to unambiguously constrain solutions to the fourth-order Jeans equation for a given mass model and (second-order) anisotropy $\beta,$ leading to a problem that is similar to the original mass-anisotropy degeneracy. One approach to this problem is to impose a prescription that fixes $\beta^{\prime}.$ This can be done, for example, by assuming distribution function models which predict these (Łokas, 2002; Łokas et al., 2005). More generally, as explored in Dejonghe & Merritt (1992); Richardson & Fairbairn (2013), assuming a particular form for the distribution function amounts to imposing a prescription to solve all the moments of the Jeans equations (beyond fourth order included). A problem with this approach, however, is that it may restrict the space of solutions unnecessarily and bias the models, unless one is able to be sufficiently judicious and general in specifying a form for the distribution function.
@@
In this way, one can exploit the increasing abundance of PM data found in nearby galactic systems and star clusters to constrain the fourth-order anisotropy, leading to a full three-dimensional and self-consistent modeling of both the second and fourth-order moments along the three observable LOS and PM components. To the knowledge of the authors, this is a new point that has not been made in the literature before. [^9] We note that Evslin & Del Popolo (2017) independently introduced fourth-order PMs to quantify departures from dynamical equilibrium. However, to the authors’ knowledge, our work is the first in which these are explicitly solved for from the Jeans equations and used to break the mass-anisotropy degeneracy, and is therefore a novelty in this respect.

### v2

@@ 2.1 An extended Jeans analysis up to fourth order
is the ‘velocity anisotropy’, [^6] Unlike in e.g. 2025A&A...693A.104B, $\sigma_{t}$ corresponds to the total 2D tangential velocity dispersion, rather than the 1D averaged one (thus the factor of two). This is purely notational, but we have adopted this change because we found it more convenient to work with this quantity for the subsequent parts of this section. and we assume Newtonian weak-field gravity with an enclosed mass $M(r)$ within the 3D radial coordinate $r$.
As discussed in 2025A&A...693A.104B, the assumption of non-rotation can be relaxed to some extent by solving for the second velocity moment directly, which we do. This is done by subtracting only global averages to the velocity distribution, leaving potential rotation signatures present. These are then reabsorbed when solving the Jeans equation, leading to a quadrature term that is implicitly added to the pure velocity dispersion (see e.g. the discussion in 2016ApJ...821...44F).
The solution for the radial velocity dispersion is then given by (2005MNRAS.363..705M; 1994MNRAS.270..271V):
@@
By considering these three separate components, provided that one has enough data, it is possible to alleviate the well-known problem of the degeneracy between the cumulative mass, $M(r)$, and stellar velocity anisotropy, $\beta(r)$ (1990AJ.....99.1548M; Wilkinson:2001ut; 2003MNRAS.343..401L; 2009MNRAS.395...76D; 2017MNRAS.471.4541R). Further, one can resort to higher moments of the velocity distribution to place additional constraints that have also been found to alleviate this problem. One popular choice is to use VSPs (1990AJ.....99.1548M; 2014MNRAS.441.1584R; 2017MNRAS.471.4541R), which are two numbers obtained by integrating the luminosity-weighted, fourth-order LOS velocity moment to infinity, thereby marginalizing out any higher-order velocity anisotropy. This approach is used in GravSphere1 and GravSphere1.5.
The key advantage of using VSPs over explicit modeling of the higher-order moment equations is that no higher-order analog of $\beta(r)$ is required. However, a key disadvantage is that the VSPs require integrals to infinity over the surface brightness profile and fourth velocity moment, which may both be poorly constrained at large radii. Furthermore, their large radius behavior could be adversely affected by disequilibrium due to tides. Both effects can lead to bias (e.g. 2017MNRAS.471.4541R; 2018MNRAS.481..860R; 2024MNRAS.535.1015D; 2025arXiv250418617T). For this reason, in GravSphere2 we take a different approach that – as we shall show (Sec. 4) – is less biased. Instead of using VSPs, we explicitly solve the fourth-order Jeans equations, introducing a flexible functional form for the associated higher-order anisotropy that we can then marginalize over. This is more similar to the approach introduced in 2013MNRAS.432.3361R. Explicitly, we solve the fourth-order moment Jeans equations (1990AJ.....99.1548M):
@@
As shown in 2013MNRAS.432.3361R, this system of equations can be further simplified, without loss of generality, by first introducing the fourth-order anisotropy analog:
@@
As for the second-order counterparts, we now seek to obtain the projected quantities which are directly related to observations. To express these purely in terms of $\braket{v^{4}_{r}}$, following 2013MNRAS.432.3361R, first, it follows from Eq. (12) that:
@@
As noted in 2013MNRAS.432.3361R, this approach constitutes a significant theoretical improvement over second-order Jeans methods by ensuring physical consistency of a given mass model with higher-order velocity moments. This is not trivially warranted, as Jeans models can in some cases lead to unphysical solutions implying negative distribution functions, or be inconsistent with the velocity distribution that is implicitly assumed to map velocity dispersions (1992ApJ...391..531D; 2013MNRAS.432.3361R) (this is explicitly addressed in the Sec. 2.2). It also has the promise of serving as a means to further constrain mass models by considering additional higher-order moments (2013MNRAS.432.3361R).
One potential disadvantage, however, which is also noted in 2013MNRAS.432.3361R, is that the constraining power of this approach can be compromised due to the fact that a fully general treatment leads to the introduction of the fourth-order anisotropy analog $\beta^{\prime},$ meaning that one may not be able to unambiguously constrain solutions to the fourth-order Jeans equation for a given mass model and (second-order) anisotropy $\beta,$ leading to a problem that is similar to the original mass-anisotropy degeneracy. One approach to this problem is to impose a prescription that fixes $\beta^{\prime}.$ This can be done, for example, by assuming distribution function models which predict these (2002MNRAS.333..697L; 2005MNRAS.363..918L). More generally, as explored in 1992ApJ...391..531D; 2013MNRAS.432.3361R, assuming a particular form for the distribution function amounts to imposing a prescription to solve all the moments of the Jeans equations (beyond fourth order included). A problem with this approach, however, is that it may restrict the space of solutions unnecessarily and bias the models, unless one is able to be sufficiently judicious and general in specifying a form for the distribution function.
@@
In this way, one can exploit the increasing abundance of PM data found in nearby galactic systems and star clusters to constrain the fourth-order anisotropy, leading to a full three-dimensional and self-consistent modeling of both the second and fourth-order moments along the three observable LOS and PM components. To the knowledge of the authors, this is a new point that has not been made in the literature before. [^8] We note that 2017ApJ...841...90E independently introduced fourth-order PMs to quantify departures from dynamical equilibrium. However, to the authors’ knowledge, our work is the first in which these are explicitly solved for from the Jeans equations and used to break the mass-anisotropy degeneracy, and is therefore a novelty in this respect.